# Essential difference between 2D and 3D from the perspective of real-space renormalization group

### v2

@@ Abstract
We point out that area laws of quantum-information concepts indicate limitations of block transformations as well-behaved real-space renormalization group (RG) maps, which in turn guides the design of better RG schemes. Mutual-information area laws imply the difficulty of Kadanoff’s block-spin method in two dimensions (2D) or higher due to the growth of short-scale correlations among the spins on the boundary of a block. A leap to the tensor-network RG, in hindsight, follows the guidance of mutual information and is efficient in 2D, thanks to its mixture of quantum and classical perspectives and the saturation of entanglement entropy in 2D. In three dimensions (3D), however, entanglement grows according to the area law, posing a threat to 3D block-tensor map as an apt RG transformation. As a numerical evidence, we show that estimations of 3D Ising critical exponents fail to improve by retaining more couplings. As a guidance to proceed, a tensor-network toy model is proposed to capture the 3D entanglement-entropy area law.
Introduction.— The renormalization group (RG) idea is an overarching theoretical framework for both elementary particle physics and statistical physics of complex systems [1, 2]. It provides a modern view of the regularization process of removing infinities in quantum fields of elementary particles; and the concepts of RG relevant and irrelevant directions have deepened our understanding of universality and emergent phenomena ubiquitous in complex systems. The essence of the RG idea is the separation of universal physics from microscopic details using a divide-and-conquer strategy. A judicious cleanse of microscopic details is the key for designing systematically-improvable RG techniques because the computational power can be saved for universal physics. Wilson’s momentum-shell RG combined with $\epsilon$-expansion [3] is the first demonstration of the power of this separation of scales, organized according to the Fourier modes of Landau-Ginzburg-Wilson field. However, for non-perturbative approaches, like real-space RG, the design of a good RG transformation is more of an art than a science [4].
@@
How do area laws have anything to do with real-space RG transformations? The connection is due to the fact that a block transformation, as the starting point of many real-space RG, generates increasingly larger blocks of the original system, which encapsulates “information” scaling according to area laws. A well-known interplay between area laws and RG is between entanglement scaling of the ground state of an (1+1)D gapped quantum system and the success of density matrix renormalization group (DMRG) [5]. Entanglement in a ground state $|0\rangle$ can be quantified in the following way: divide the system into two parts: a block $B$ and its complement $\bar{B}$; define the density matrix of $B$ by tracing out the Hilbert space of $\bar{B}$ as
@@
An area law means a scaling behavior like the boundary area $|\partial B|$ of the block: $S(\rho_{B})=O(|\partial B|)$ in the case of entanglement entropy [6]. In a (1+1)D gapped system, the boundaries are two points and entanglement entropy becomes a constant for a block $B$ with a size larger than the correlation length; this saturation of entanglement entropy justifies the DMRG approximation of keeping only a constant number of states of the block [5].
The focus here is on the properties of real-space RG in the context of classical statistical mechanics, specifically about block transformations, using area laws as a tool. One hopes that a better understanding of the nature and limitations of such a basic operation can guide the design of efficient and well-behaved real-space RG transformations.
Properties of an apt RG.—Before assessing a real-space RG, it is better to have a clear-cut answer to the first question: What is a good RG transformation in a statistical physics setting? We list three basic properties:
[Approximate $Z$ condition]: The partition function $Z$ after an RG map should be as close to the original one as possible.
@@
[UV-free condition]: An RG map should strive to identify and clean short-scale physics as much as possible, so the finite computation power and memory are saved for the universal properties.
The first [Approximate $Z$ condition] is the defining property of a RG transformation in statistical physics. Suppose we have the reduced energy of a system as $\mathcal{H}_{\{K_{j}\}}[\{s_{\mathbf{r}}\}]$, where $s_{\mathbf{r}}$ denotes a spin located at lattice site $\mathbf{r}$ and $K_{j}$ is the $j$-th coupling constant parametrizing the energy function $\mathcal{H}$. The statistical weight of a given configuration of spins is the Boltzmann factor $w(\{K_{j}\};\{s_{\mathbf{r}}\})=e^{-\mathcal{H}_{\{K_{j}\}}[\{s_{\mathbf{r}}\}]}$. The partition function $Z\left(\{K_{j}\}\right)=\sum_{\{s_{\mathbf{r}}\}}w(\{K_{j}\};\{s_{\mathbf{r}}\})$, as a function of coupling constants, generates various correlation functions. A real-space RG transformation is a rule of mapping the original spins $\{s_{\mathbf{r}}\}$ to renormalized ones $\{s^{\prime}_{\mathbf{r}^{\prime}}\}$, inducing a map in coupling-constant space $\mathcal{T}:\{K_{j}\}\mapsto\{K_{j}^{\prime}\}$, such that the new system with energy function $\mathcal{H}_{\{K_{j}^{\prime}\}}[\{s^{\prime}_{\mathbf{r}^{\prime}}\}]$ has the same partition function as the old one $Z\left(\{K_{j}^{\prime}\}\right)=Z\left(\{K_{j}\}\right)$. A direct consequence of [Approximate $Z$ condition] is that all correlation functions between renormalized spins in larger scales are encoded in $Z\left(\{K_{j}^{\prime}\}\right)$ of a coarser system without knowing anything about the original one [7].
The second property, [RG flow condition], makes it possible to identify different phases of matter from different isolated fixed points under the RG map $\mathcal{T}:\{K_{j}^{*}\}\mapsto\{K_{j}^{*}\}$; universal properties like critical exponents can then be extracted from the linearized RG transformation at a critical fixed point. One may wonder whether [RG flow condition] is related to the defining property [Approximate $Z$ condition]. In practice, truncation is usually necessary. It is possible that an unwise truncation scheme causes incorrect RG flows. However, we want to point out that truncation is not the only cause of incorrect RG flows. In fact, recently, it is demonstrated that a block transformation, even if performed exactly, leads to unsatisfactory RG flows [8, 9].
@@
Kadanoff’s block spin.—We go back all the way to the origin of real-space RG and examine Kadanoff’s block-spin idea [10]. A block-spin method, like decimation, if performed exactly, entails an indefinite growth of couplings. Can the usual practice of retaining a constant number of couplings be justified? Mutual information and its area law for a system in thermal equilibrium shed some light on this question [11, 12]. Intuitively, the correlation between the block of spins $B$ and its environment $E$ is mediated through a strip of spins around their interface, roughly with the width of a characteristic length scale $\xi$ of the system. This correlation can be quantified using mutual information $I(B,E)$ and is bounded by the number of spins inside the strip, $I(B,E)\leq 4|\partial B|\xi$, where $|\partial B|$ is the area of the boundary of the block. In the worst case, for a classical statistical system in $d\geq 2$, the correlations between the block and its environment increase according to $L^{d-1}$ as the size $L$ of the block grows under RG. The tension between the growth of mutual information and the constant number of couplings retained during RG suggests one possible explanation of why many real-space RG schemes based on spin pictures fail to produce systematically improvable estimation of critical exponents [13, 14].
@@
Block tensor in 2D.—The two essential ideas of tensor-network renormalization group (TNRG) are
(i) focusing on spins at the boundary of a block and
(ii) a mixture of quantum and classical view, so we can use unitary transformations to identify important states.
For a large block, the partition function with boundary spin fixed while inner spins summed over represents the ground state $|g\rangle$ of a quantum Hamiltonian that corresponds to the transfer matrix acting on the boundary spins [14, 15, 16]. The advantage of the quantum perspective is that area laws of entanglement for quantum systems become applicable to classical statistical systems, with two reductions of dimensionality—one from classical-to-quantum mapping and the other from an area law.
Specifically, imagine a 2D classical system with partition function $Z$ and divide it into square blocks. Each block, as a $4$-leg tensor, represents the ground state wavefunction [15] of spins located at the boundary,
@@
where, for example, $\vec{s}_{l}$ denotes all spins on the left edge of the square. With copies of the block as tiles, the partition function $Z$ can be represented by the contraction $Z_{N}(A)$ of a tensor network of $N$ copies of $A$, so $Z=Z_{N}(A)$ [15]. The 2D block-tensor transformation with rescaling factor $b$ has three steps—(i) grouping a larger square block of $b^{2}$ copies of tensor $A$, (ii) summing over spins on inner edges, and (iii) a unitary transformation on spins sitting on each outer edge of the larger square block to identify important states according to their weight in the entanglement spectrum of the larger square block. The Hilbert space corresponding to the larger square block is similarly divided into $4$ groups according to $4$ edges of the larger square block, which defines a coarse tensor $A^{\prime}$ (see Fig. 1 for a 3D block tensor). This defines a map in tensor space,
@@
We have $Z_{N}(A)=Z_{N/b^{2}}(A^{\prime})$ by construction, provided no truncation made in Step (iii). According to the entanglement area law of (1+1)D gapped systems, the density matrix $\rho_{\vec{s}_{l}}$ of the subsystem $\vec{s}_{l}$ has entanglement entropy $S(\rho_{\vec{s}_{l}})\to S_{0}\sim\log\xi$ if the statistical system has characteristic length scale $\xi$ and the block size $L\gg\xi$ [5, 17]. As a result, as long as the number of states $\chi$ retained in Step (iii) is $O(\xi)$, the approximation is quasi-exact [15]. We see that the role of entanglement entropy in TNRG is twofold—it implies how well the partition function is preserved and also serves as a single-number characterization of the tensor, reflecting [Approximate $Z$ condition] and [RG flow condition] respectively.
@@
The above arguments imply that the 2D block tensor can be an apt RG transformation. [Approximate $Z$ condition] is well satisfied due to the saturation of entanglement entropy. This is a great step forward from Kadanoff’s block-spin method. This advance is achieved thanks to the aid of density matrix $\rho_{\vec{s}_{l}}$ for the identification of short-scale correlations among spins $\vec{s}_{l}$ on an edge of a square block, a manifest improvement on [UV-free condition]. However, $\rho_{\vec{s}_{l}}$ still preserves some short-scale correlations between the spin group $\vec{s}_{l}$ and the rest $\vec{s}_{u}\cup\vec{s}_{d}\cup\vec{s}_{r}$, which is reflected in $S_{0}\sim\xi$. The remnant short-scale entanglement is responsible for continua of fixed-point tensors in the neighborhood of trivial phases [15, 18], which was noticed since the birth of 2D TNRG methods—[RG flow condition] is violated. Efforts to filter the remnant entanglement lead to beyond-simple-block 2D TNRG schemes with a full satisfaction of all three properties, as numerical methods [18, 19, 20, 21, 22, 23, 24] and as exact RG transformations [8, 9].

### v3

@@ Abstract
We point out that area laws of quantum-information concepts indicate limitations of block transformations as well-behaved real-space renormalization group (RG) maps, which in turn guides the design of better RG schemes. Mutual-information area laws imply the difficulty of Kadanoff’s block-spin method in two dimensions (2D) or higher due to the growth of short-scale correlations among the spins on the boundary of a block. A leap to the tensor-network RG, in hindsight, follows the guidance of mutual information and is efficient in 2D, thanks to its mixture of quantum and classical perspectives and the saturation of entanglement entropy in 2D. In three dimensions (3D), however, entanglement grows according to the area law, posing a threat to a 3D block-tensor map as an apt RG transformation. As numerical evidence, we show that estimates of 3D Ising critical exponents fail to converge with respect to the RG step, making the 3D block-tensor map an unreliable RG method. Moreover, the estimates do not improve as more coupling constants are retained. As a guidance to proceed, a tensor-network toy model is proposed to capture the 3D entanglement-entropy area law.
keywords:
Real space renormalization group, Tensor Network, Entanglement entropy
1 Introduction
The renormalization group (RG) idea is an overarching theoretical framework for both elementary particle physics and statistical physics of complex systems [wilson-kogut1974, wilson1975kondo]. It provides a modern view of the regularization process of removing infinities in quantum fields of elementary particles; and the concepts of RG relevant and irrelevant directions have deepened our understanding of universality and emergent phenomena ubiquitous in complex systems. The essence of the RG idea is the separation of universal physics from microscopic details using a divide-and-conquer strategy. A judicious cleanse of microscopic details is the key for designing systematically-improvable RG techniques because the computational power can be saved for universal physics. Wilson’s momentum-shell RG combined with $\epsilon$-expansion [wilson-fisher1972] is the first demonstration of the power of this separation of scales, organized according to the Fourier modes of Landau-Ginzburg-Wilson field. However, for non-perturbative approaches, like real-space RG, the design of a good RG transformation is more of an art than a science [fisher1998].
@@ 1 Introduction
How do area laws have anything to do with real-space RG transformations? The connection is due to the fact that a block transformation, as the starting point of many real-space RG, generates increasingly larger blocks of the original system, which encapsulates “information” scaling according to area laws. An example of how the entanglement-entropy area law can guide the design of an efficient RG map for a quantum chain is Entanglement Renormalization invented by Vidal [vidal2003, Vidal:2007, Vidal:2010:ER]. Entanglement in a ground state $|0\rangle$ can be quantified in the following way: divide the system into two parts: a block $B$ and its complement $\bar{B}$; define the density matrix of $B$ by tracing out the Hilbert space of $\bar{B}$ as
@@
An area law means a scaling behavior like the boundary area $|\partial B|$ of the block: $S(\rho_{B})=O(|\partial B|)$ in the case of entanglement entropy [Eisert2010area]. In a (1+1)D gapped system, the boundaries are two points and entanglement entropy becomes a constant for a block $B$ with a size larger than the correlation length; this saturation of entanglement entropy justifies the DMRG approximation of keeping only a constant number of states of the block [vidal2003].
We emphasize here that the two parts of the system are divided in real space. The amount of entanglement depends drastically on the space where the state is represented. For example, a non-interacting fermionic system can be completely factorized in momentum space, but obeys an area law in real space, sometimes with a logarithmic correction [Wolf:2006]. One can try circumventing the area law in real space by analyzing models expressed in momentum space, like lowest Landau level orbits on a two-sphere [FuzzySphere:2023].
The focus here is on the properties of real-space RG in the context of classical statistical mechanics, specifically about block transformations, using area laws as a tool. One hopes that a better understanding of the nature and limitations of such a basic operation can guide the design of efficient and well-behaved real-space RG transformations. We restrict our discussion to these real-space RG transformations where the RG map is constructed directly by imposing the condition that the partition function should be well approximated. Examples of such kind include Migdal-Kadanoff (MK) bond-moving approach [Migdal:1975a, Kadanoff:1976], Wilson’s numerical implementation of a block-spin map [wilson1975kondo] and other variations of MK scheme [Kadanoff:1975:var, Martinelli:Parisi:1981]. The modern tensor-network renormalization group (TNRG) methods [levin-nave] also belong to this type; they can be seen as a generalization of the MK approach [Meurice:2013]. In this sense, the Monte Carlo RG [Swendsen:1979] does not fit in the scope of this paper. The reason is that the Monte Carlo RG involves a stochastic sampling process; it complicates the discussion here, where the emphasis is the properties of an RG map per se.
The argument in this paper is inspired by Levin and Nave [levin-nave]. When implementing the first block transformation in tensor-network language, they use an area-law argument to show that the numerical truncations in their scheme cannot be guaranteed to be accurate in 2D critical and 3D systems. They see the TNRG as a numerical method for approximating the partition function. We expand their argument to show further implications of the area law for understanding the properties of TNRG in the context of the Wilsonian RG framework. Our argument helps clarify why the difficulty of the TNRG in 3D is qualitatively different from 2D, as well as some confusions in the 3D TNRG calculations [hotrg2012, wang2014, jha-3dpotts].
2 Properties of an apt RG
Before assessing a real-space RG, it is better to have a clear-cut answer to the first question: What is a good RG transformation in a statistical physics setting? We list three basic properties:
[Approximate $Z$ condition]: The partition function $Z$ after an RG map should be close to the original one.
@@ 2 Properties of an apt RG
[UV-free condition]: An RG map should strive to identify and filter out short-scale physics as much as possible, so the finite computation power and memory are saved for the universal properties.
The first [Approximate $Z$ condition] is the defining property of a RG transformation in statistical physics. Suppose we have the reduced energy of a system as $\mathcal{H}_{\{K_{j}\}}[\{s_{\mathbf{r}}\}]$, where $s_{\mathbf{r}}$ denotes a spin located at lattice site $\mathbf{r}$ and $K_{j}$ is the $j$-th coupling constant parametrizing the energy function $\mathcal{H}$. The statistical weight of a given configuration of spins is the Boltzmann factor $w(\{K_{j}\};\{s_{\mathbf{r}}\})=e^{-\mathcal{H}_{\{K_{j}\}}[\{s_{\mathbf{r}}\}]}$. The partition function $Z\left(\{K_{j}\}\right)=\sum_{\{s_{\mathbf{r}}\}}w(\{K_{j}\};\{s_{\mathbf{r}}\})$, as a function of coupling constants, generates various correlation functions. A real-space RG transformation is a rule of mapping the original spins $\{s_{\mathbf{r}}\}$ to renormalized ones $\{s^{\prime}_{\mathbf{r}^{\prime}}\}$, inducing a map in coupling-constant space $\mathcal{T}:\{K_{j}\}\mapsto\{K_{j}^{\prime}\}$, such that the new system with energy function $\mathcal{H}_{\{K_{j}^{\prime}\}}[\{s^{\prime}_{\mathbf{r}^{\prime}}\}]$ has the same partition function as the old one $Z\left(\{K_{j}^{\prime}\}\right)=Z\left(\{K_{j}\}\right)$. In a practical calculation, the partition function in the new description should be a good approximation of the original one. A direct consequence of [Approximate $Z$ condition] is that all correlation functions between renormalized spins in larger scales are encoded in $Z\left(\{K_{j}^{\prime}\}\right)$ of a coarser system without knowing anything about the original one [cardy-book].
The second property, [RG flow condition], makes it possible to identify different phases of matter from different isolated fixed points under the RG map $\mathcal{T}:\{K_{j}^{*}\}\mapsto\{K_{j}^{*}\}$; universal properties like critical exponents can then be extracted from the linearized RG transformation at a critical fixed point. One may wonder whether [RG flow condition] is related to the defining property [Approximate $Z$ condition]. In practice, truncation is usually necessary. It is possible that an unwise truncation scheme causes incorrect RG flows. However, we want to point out that truncation is not the only cause of incorrect RG flows. In fact, Kadanoff [Kadanoff:1976] argued that the spin decimation, even if performed exactly, leads to a spin-spin correlation that does not change under a change of length scale; this indicates that the spin decimation always produce fixed points with the exponent $\beta/\nu=\Delta_{\sigma}=0$. Recently, it is demonstrated that an exact block transformation in tensor-network language also gives unsatisfactory RG flows [kennedy-hiT, kennedy-lowT], where both the high- and the low-temperature fixed points form a continuum, instead of isolated fixed points.
@@
3 Kadanoff’s block spin
We go back all the way to the origin of real-space RG and examine Kadanoff’s block-spin idea [kadanoff-spinblock]. A block-spin method, like decimation, if performed exactly, entails an indefinite growth of couplings. Can the usual practice of retaining a constant number of couplings be justified? Mutual information and its area law for a system in thermal equilibrium shed some light on this question [verstraete2008, verstraete-talk]. Intuitively, the correlation between the block of spins $B$ and its environment $E$ is mediated through a strip of spins around their interface, roughly with the width of a characteristic length scale $\xi$ of the system [verstraete2008]. This correlation can be quantified using mutual information $I(B,E)$ and is bounded by the number of spins inside the strip, $I(B,E)\leq 4|\partial B|\xi$, where $|\partial B|$ is the area of the boundary of the block. In the worst case, for a classical statistical system in $d\geq 2$, the correlations between the block and its environment increase according to $L^{d-1}$ as the size $L$ of the block grows under RG. The tension between the growth of mutual information and the constant number of couplings retained during RG suggests one possible explanation of why many real-space RG schemes based on spin pictures fail to produce systematically improvable estimates of critical exponents [kardar-book, shankar-qftbook].
@@ 3 Kadanoff’s block spin
4 Block tensor and entanglement entropy
The two essential ideas of tensor-network renormalization group (TNRG) are [levin-nave]
(1) focusing on spins at the boundary of a block and
(2) a mixture of quantum and classical view, so we can use unitary transformations to identify important states.
For a large block, the partition function with boundary spin fixed while inner spins summed over represents the ground state $|g\rangle$ of a quantum Hamiltonian that corresponds to the transfer matrix acting on the boundary spins [shankar-qftbook, levin-nave, kogut1979]. The advantage of the quantum perspective is that area laws of entanglement for quantum systems become applicable to classical statistical systems, with two reductions of dimensionality—one from classical-to-quantum mapping and the other from an area law. One salient feature of the TNRG is that scaling dimensions of a variety of scaling operators are contained in the linearized tensor RG map, if the linearization is performed in a particular manner [lyu-kawashima2021, Ebel:2024LDO]. For example, in Fig. 5, the linearized RG map contains scaling dimensions of the energy-momentum tensor $T_{ij}$ and various descendants like $\partial_{i}\epsilon,\partial_{i}\sigma$ with $i=x,y,z$ of the 3D Ising criticality, most of them with an approximate degeneracy anticipated by conformal field theory. This has not been demonstrated in previous RG methods.
Specifically, imagine a 2D classical system with partition function $Z$ and divide it into square blocks. Each block, as a $4$-leg tensor, represents the ground state wavefunction [levin-nave] of spins located at the boundary,
@@ 4 Block tensor and entanglement entropy
where, for example, $\vec{s}_{l}$ denotes all spins on the left edge of the square. With copies of the block as tiles, the partition function $Z$ can be represented by the contraction $Z_{N}(A)$ of a tensor network of $N$ copies of $A$, so $Z=Z_{N}(A)$ [levin-nave]. The density matrix of a leg of the tensor $A$ is obtained according to Eq. (1) by choosing a leg as the system $B$ while the remaining legs as $\bar{B}$. For example, if one chooses leg $\vec{s}_{l}$, the density matrix is [^1] Notice that the complex conjugate is not needed for classical statistical models since the tensor is real-valued.
$\rho_{\vec{s}_{l}^{\prime}\vec{s}_{l}}=\sum_{\vec{s}_{r},\vec{s}_{u},\vec{s}_{d}}A_{\vec{s}_{l}^{\prime},\vec{s}_{r},\vec{s}_{u},\vec{s}_{d}}A_{\vec{s}_{l},\vec{s}_{r},\vec{s}_{u},\vec{s}_{d}},$ (4)
from which the entanglement entropy can be calculated according to Eq. (2). This single-tensor-level entanglement entropy serves as a single-number characterization of the tensor, which can help understand [RG flow condition] of a block-tensor map. Similarly, a density matrix can be obtained for a larger patch of the tensor network.
Figure 1: Block-tensor transformation and its spin picture with $b=2$. The “Block” step in this Figure is the Step (i), while the “Tensor contraction” step corresponds to Step (ii) and (iii). The triangular tensor in (a) is the unitary transformation that identifies important states. In numerical calculations, truncations of the state are according to the eigenvalue spectrum of the density matrix of the larger square block after the “Block” step. When truncations happen, the triangular tensors are known as isometric tensors [tnr2015].
The 2D block-tensor transformation with rescaling factor $b$ has three steps (Fig. 1 shows a $b=2$ transformation)—
Step (i): grouping a larger square block of $b^{2}$ copies of tensor $A$,
Step (ii): summing over spins on inner edges, and
Step (iii): a unitary transformation on spins sitting on each outer edge of the larger square block to identify important states according to the eigenvalue spectrum of the density matrix of the larger square block (see Fig. 1).
These three steps are general in higher dimensions. In 2D, the Hilbert space corresponding to the larger square block is similarly divided into $4$ groups according to $4$ edges of the larger square block, which defines a coarse tensor $A^{\prime}$ This defines a map in tensor space,
@@
We have $Z_{N}(A)=Z_{N/b^{2}}(A^{\prime})$ by construction, provided no truncation made in Step (iii). From Fig. 1, we see that under a block-tensor transformation, the size of the spin block that the coarse-grained tensor $A^{\prime}$ represents becomes $b$ times as large as what $A$ represents.
In numerical calculations, usually the first $\chi$ eigenstates of the density matrix of the larger square block with the $\chi$ largest eigenvalues are chosen, the remaining states being truncated. The error of this truncation can be approximated by the $\chi$-th eigenvalue if the density matrix is normalized to have unit trace. Recall that a larger von Neumann entropy of an eigenvalue spectrum usually corresponds to a slow decay of the spectrum, thus indicating a larger RG truncation error. Therefore, the second role of entanglement entropy is how well the partition function is preserved under the RG transformation, reflecting [Approximate $Z$ condition] of a block-tensor map.
In summary, the role of entanglement entropy in the block-tensor map is twofold—it implies how well the partition function is preserved and also serves as a single-number characterization of the tensor, reflecting [Approximate $Z$ condition] and [RG flow condition] respectively. A larger entanglement entropy indicates a larger RG truncation error, thus a poor RG approximation for the partition function. A changing of the entanglement entropy with respect to the size of the system, or the RG step, indicates the tensor itself is not fixed under the RG transformation. These two roles of entanglement entropy will be employed to understand the properties of the block-tensor map in 2D and 3D in the following section.
5 Is the block-tensor map an apt RG?
5.1 In 2D
@@
According to the entanglement area law of (1+1)D gapped systems, the density matrix $\rho_{\vec{s}_{l}}$ of the subsystem $\vec{s}_{l}$ has entanglement entropy $S(\rho_{\vec{s}_{l}})\to S_{0}\sim\log\xi$ if the statistical system has characteristic length scale $\xi$ and the block size $L\gg\xi$ [vidal2003, cardy-eeqft-intro]. As a result, as long as the number of states $\chi$ retained in Step (iii) is $O(\xi)$, the approximation is quasi-exact [levin-nave].
@@
The above arguments imply that the 2D block tensor can be an apt RG transformation. [Approximate $Z$ condition] is well satisfied due to the saturation of entanglement entropy. This is a great step forward from Kadanoff’s block-spin method. This advance is achieved thanks to the aid of density matrix $\rho_{\vec{s}_{l}}$ for the identification of short-scale correlations among spins $\vec{s}_{l}$ on an edge of a square block, a manifest improvement on [UV-free condition]. However, $\rho_{\vec{s}_{l}}$ still preserves some short-scale correlations between the spin group $\vec{s}_{l}$ and the rest $\vec{s}_{u}\cup\vec{s}_{d}\cup\vec{s}_{r}$, which is reflected in $S_{0}\sim\xi$. The remnant short-scale entanglement is responsible for continua of fixed-point tensors in the neighborhood of trivial phases [levin-nave, gu-wen2009], which was noticed since the birth of 2D TNRG methods—[RG flow condition] is violated. Efforts to filter the remnant entanglement lead to beyond-simple-block 2D TNRG schemes with a full satisfaction of all three properties, as numerical methods [gu-wen2009, tnr2015, tnr-plus2017, loop-tnr2017, gilt2018, harada2018, fet2018] and as exact RG transformations [kennedy-hiT, kennedy-lowT].
5.2 Exact 3D block-tensor map